An undersea thermal vent emits ocean water which was superheated by a magma chamber below the ocean floor. The temperature of the superheated water is 400° C. A submersible sends a steel probe of volume 0.0008 m3 through the heated water from the surrounding cold ocean t 5° C. By how much does the volume of the steel probe change? Take steel = 36 x 10-6 (°C)-1.
1
Expert's answer
2020-04-28T09:48:08-0400
"\Delta V=\beta V\Delta T"

"\Delta V=(36\cdot 10^{-6} )(0.0008)(400-5)=1.1\cdot 10^{-5}\ m^3"